In 20 years, Lester will be 5 times his age 8 years ago. How old is he at present?
Answer by ikleyn(52794)

.

    x + 20 = 5*(x-8)


It is your basic equation, from the condition.


Solve it


    x + 20  = 5x - 40

    20 + 40 = 5x - x

       60   = 4x

    x = 60/4 = 15.


ANSWER.  Lester is 15 years old.

Solved.